# High-dimensional Gaussian and bootstrap approximations for robust means

### v3

@@ Appendix D Proofs for Section 3
$B_{n}\mathrel{\mathop{\ordinarycolon}}=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}-\mathbb{P}\mathinner{\bigl(\hat{Z}\in H\mid\tilde{X}_{1},\ldots,\tilde{X}_{n}\bigr)}\Bigr\rvert}$.
First, $\rho_{n,W}\leq\mathfrak{A}_{n}$, where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in Theorem 2.1. Next, by the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 of Chernozhuokov et al. (2022) (cf. also Proposition 2 in Chernozhukov et al. (2023a)),
@@
$B_{n}\leq C\mathinner{\Bigl(\log^{2}(d)\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\hat{\Sigma}_{n,j,k,\mathrm{PSD}}-\Sigma_{j,k}\bigr\rvert}\Bigr)}^{1/2}\leq C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\Biggr)}^{1/2},$
the last inequality holding with probability at least $1-\frac{12}{n}$ by Remark D.1 and $C$ being a constant depending on $b_{1},b_{2},c$ and $m$ only. That $\hat{\rho}_{n,W}\to 0$ in probability follows from from (B.7)–(B.9) of Lemma B.4. ∎
We now prove Theorem 3.1 by showing that $\tilde{\Sigma}_{n}$ is suitably close to $\hat{\Sigma}_{n}$.
@@
The theorem is proved by showing that for a constant $C$ depending only on $b_{2},c$, and $m$,
$\mathbb{P}\mathinner{\biggl(\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\hat{\Sigma}_{n,j,k}\bigr\rvert}>C\mathinner{\Bigl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{\frac{1}{2}-\frac{1}{m}}\Bigr)}\biggr)}\leq\frac{12}{n},$ (D.5)
which together with the triangle inequality and Theorem D.1 yields the desired conclusion. To prove (D.5), recall the notation prior to (D.1), fix $1\leq j,k\leq d$, and note that with
$A_{n,j,k}\mathrel{\mathop{\ordinarycolon}}=\mathinner{\Bigl\{i\in\mathinner{\bigl\{1,\ldots,N\bigr\}}\mathrel{\mathop{\ordinarycolon}}\hat{\alpha}_{j,k}\leq\tilde{U}_{i,j,k}\leq\hat{\beta}_{j,k}\text{ and }\hat{a}_{j}\leq\tilde{Y}_{i,j}\leq\hat{b}_{j}\text{ and }\hat{a}_{k}\leq\tilde{Y}_{i,k}\leq\hat{b}_{k}\Bigr\}},$
one has that
$\phi_{\hat{\alpha}_{j,k},\hat{\beta}_{j,k}}\mathinner{(\tilde{U}_{i,j,k})}-\phi_{\hat{a}_{j},\hat{b}_{j}}(\tilde{Y}_{i,j})\phi_{\hat{a}_{k},\hat{b}_{k}}(\tilde{Y}_{i,k})=\tilde{U}_{i,j,k}-\tilde{Y}_{i,j}\tilde{Y}_{i,k}=0\qquad\text{for }i\in A_{n,j,k}.$
Hence,
$\tilde{\Sigma}_{n,j,k}-\hat{\Sigma}_{n,j,k}=\frac{1}{N}\sum_{i\in A_{n,j,k}^{c}}\mathinner{\bigl[\phi_{\hat{\alpha}_{j,k},\hat{\beta}_{j,k}}\mathinner{(\tilde{U}_{i,j,k})}-\phi_{\hat{a}_{j},\hat{b}_{j}}(\tilde{Y}_{i,j})\phi_{\hat{a}_{k},\hat{a}_{k}}(\tilde{Y}_{i,k})\bigr]},$
and for every $i\in A_{n,j,k}^{c}$
$\mathinner{\!\Bigl\lvert\phi_{\hat{\alpha}_{j,k},\hat{\beta}_{j,k}}\mathinner{(\tilde{U}_{i,j,k})}-\phi_{\hat{a}_{j},\hat{b}_{j}}(\tilde{Y}_{i,j})\phi_{\hat{a}_{k},\hat{b}_{k}}(\tilde{Y}_{i,k})\Bigr\rvert}\leq\mathinner{\bigl(|\hat{\alpha}_{j,k}|\vee|\hat{\beta}_{j,k}|\bigr)}+\mathinner{\bigl(|\hat{a}_{j}|\vee|\hat{b}_{j}|\bigr)}\mathinner{\bigl(|\hat{a}_{k}|\vee|\hat{b}_{k}|\bigr)}.$ (D.6)
To bound the right-hand side of the previous display, observe that since $n=2N$,
$\mathinner{\!\bigl\lvert\mathinner{\bigl\{i\in\mathinner{\{1,\ldots,N\}}\mathrel{\mathop{\ordinarycolon}}\tilde{U}_{i,j,k}\neq U_{i,j,k}\bigr\}}\bigr\rvert}\leq 2\overline{\eta}_{n}N\quad\text{and}\quad\mathinner{\!\bigl\lvert\mathinner{\bigl\{i\in\mathinner{\{1,\ldots,N\}}\mathrel{\mathop{\ordinarycolon}}\tilde{Y}_{i,j}\neq Y_{i,j}\bigr\}}\bigr\rvert}\leq 2\overline{\eta}_{n}N$
[…]
$Q_{\varepsilon_{n}^{\prime}-c^{-1}\varepsilon_{n}^{\prime}}(U_{1,j,k})\leq\hat{\alpha}_{j,k}\leq\hat{\beta}_{j,k}\leq Q_{1-\varepsilon_{n}^{\prime}+c^{-1}\varepsilon_{n}^{\prime}}(U_{1,j,k})$
with probability at least $1-2\cdot\frac{\delta}{6}=1-\frac{2}{d^{2}N}$, and where the second inequality used that $\varepsilon_{n}^{\prime}\in(0,1/2)$. Therefore, with at least this probability,
$|\hat{\alpha}_{j,k}|\vee|\hat{\beta}_{j,k}|\leq|Q_{\varepsilon_{n}^{\prime}-c^{-1}\varepsilon_{n}^{\prime}}(U_{1,j,k})|\vee|Q_{1-\varepsilon_{n}^{\prime}+c^{-1}\varepsilon_{n}^{\prime}}(U_{1,j,k})|.$
$\mathinner{\bigl(|\hat{a}_{j}|\vee|\hat{b}_{j}|\bigr)}\mathinner{\bigl(|\hat{a}_{k}|\vee|\hat{b}_{k}|\bigr)}\leq$ $\mathinner{\bigl(|Q_{\varepsilon_{n}^{\prime}-c^{-1}\varepsilon_{n}^{\prime}}(Y_{1,j})|\vee|Q_{1-\varepsilon_{n}^{\prime}+c^{-1}\varepsilon_{n}^{\prime}}(Y_{1,j})|\bigr)}$
$\cdot$ $\mathinner{\bigl(|Q_{\varepsilon_{n}^{\prime}-c^{-1}\varepsilon_{n}^{\prime}}(Y_{1,k})|\vee|Q_{1-\varepsilon_{n}^{\prime}+c^{-1}\varepsilon_{n}^{\prime}}(Y_{1,k})|\bigr)}.$
$\mathinner{\bigl(|Q_{\varepsilon_{n}^{\prime}-c^{-1}\varepsilon_{n}^{\prime}}(U_{1,j,k})|\vee|Q_{1-\varepsilon_{n}^{\prime}+c^{-1}\varepsilon_{n}^{\prime}}(U_{1,j,k})|\bigr)}+$ $\mathinner{\bigl(|Q_{\varepsilon_{n}^{\prime}-c^{-1}\varepsilon_{n}^{\prime}}(Y_{1,j})|\vee|Q_{1-\varepsilon_{n}^{\prime}+c^{-1}\varepsilon_{n}^{\prime}}(Y_{1,j})|\bigr)}$
$\cdot$ $\mathinner{\bigl(|Q_{\varepsilon_{n}^{\prime}-c^{-1}\varepsilon_{n}^{\prime}}(Y_{1,k})|\vee|Q_{1-\varepsilon_{n}^{\prime}+c^{-1}\varepsilon_{n}^{\prime}}(Y_{1,k})|\bigr)}.$
Next, note that $\mathbb{E}U_{1,j,k}=\Sigma_{j,k}$$\mathbb{E}\mathinner{\lvert U_{1,j,k}-\mathbb{E}U_{1,j,k}\rvert}^{m/2}\leq 2^{m}\sigma_{m}^{m}$ (cf. (D.4)), $\mathbb{E}Y_{1,j}=0$, and $\mathbb{E}|Y_{1,j}|^{m}\leq 2^{m/2}\sigma_{m}^{m}$ for all $1\leq j,k\leq d$. Hence, Lemma G.1 implies that the previous display is bounded from above by
$\mathinner{\biggl(|\Sigma_{j,k}|+\frac{4\sigma_{m}^{2}}{(\varepsilon_{n}^{\prime}-c^{-1}\varepsilon_{n}^{\prime})^{2/m}}\biggr)}+\frac{2\sigma_{m}^{2}}{(\varepsilon_{n}^{\prime}-c^{-1}\varepsilon_{n}^{\prime})^{2/m}}\leq\frac{7\sigma_{m}^{2}}{(\varepsilon_{n}^{\prime}-c^{-1}\varepsilon_{n}^{\prime})^{2/m}},$
the inequality following from $|\Sigma_{j,k}|\leq\sigma_{2}^{2}\leq\sigma_{m}^{2}$ and $\varepsilon_{n}^{\prime}-c^{-1}\varepsilon_{n}^{\prime}\in(0,1)$. Thus, since $|A_{n,j,k}^{c}|\leq 6\varepsilon_{n}^{\prime}N$, with probability at least $1-\frac{6}{d^{2}N}$ it holds that
$\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\hat{\Sigma}_{n,j,k}\bigr\rvert}$ $\leq 6\varepsilon_{n}^{\prime}\cdot\frac{7\sigma_{m}^{2}}{(\varepsilon_{n}^{\prime}-c^{-1}\varepsilon_{n}^{\prime})^{2/m}}\leq 42\sigma_{m}^{2}[c/(c-1)]^{2/m}{\varepsilon_{n}^{\prime}}^{1-\frac{2}{m}}$
$\leq C\mathinner{\Bigl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(d^{2}N)}{N}\Bigr]}^{\frac{1}{2}-\frac{1}{m}}\Bigr)},$
where we inserted $\varepsilon_{n}^{\prime}$ from (9), used subadditivity of $z\mapsto z^{1-\frac{2}{m}}$, and $C$ is a constant depending only on $b_{2},c$, and $m$. Hence, (D.5) follows by the union bound over the $d^{2}$ entries of the covariance matrices and $N=n/2$ upon adjusting multiplicative constants. ∎
Proof of Theorem 3.2.
The proof is almost identical to that of Theorem D.2, but is included for completeness. By the triangle inequality $\tilde{\rho}_{n,W}$ is bounded from above by the sum of
@@
First, $\rho_{n,W}\leq\mathfrak{A}_{n}$, where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in Theorem 2.1. Next, by the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 of Chernozhuokov et al. (2022) (cf. also Proposition 2 in Chernozhukov et al. (2023a)),
$B_{n}\leq C\mathinner{\Bigl(\log^{2}(d)\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}\Bigr)}^{1/2}\leq C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\Biggr)}^{1/2},$
the last inequality holding with probability at least $1-\frac{24}{n}$ by Theorem 3.1 and $C$ being a constant depending on $b_{1},b_{2},c$ and $m$ only. That $\tilde{\rho}_{n,W}\to 0$ in probability follows from (B.7)–(B.9) of Lemma B.4. ∎
@@ Appendix E Proofs for Section 4
$\frac{\overline{\delta}}{\underline{\sigma}}\mathinner{\bigl(\sqrt{2\log(2d)}+2\bigr)}=\frac{\overline{\delta}}{\underline{\sigma}}\mathinner{\bigl(\sqrt{2(\log(2)+\log(d))}+2\bigr)}\leq\frac{\overline{\delta}}{\underline{\sigma}}\mathinner{\bigl(\sqrt{\log(d)}+2+\sqrt{2(\log(2)}\bigr)},$
@@
We first establish (4.1), and assume that
$C_{1}\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\leq\frac{b_{1}^{2}}{2},$ (E.3)
where $C_{1}$ is the constant from Theorem 3.1 depending only on $b_{2},c$, and $m$. This is without loss of generality, because if (E.3) does not hold, we can conclude (4.1) immediately as in the end of the proof of Theorem 2.1.
@@
$T_{n,j}=\frac{1}{\sqrt{n}\sigma_{2,j}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]},\qquad j=1,\ldots,n,$
@@
$\max_{j=1,\ldots,d}\mathinner{\!\biggl\lvert\frac{1}{\tilde{\sigma}_{n,j}}-\frac{1}{\sigma_{2,j}}\biggr\rvert}=\max_{j=1,\ldots,d}\frac{|\tilde{\sigma}_{n,j}-\sigma_{2,j}|}{\tilde{\sigma}_{n,j}\sigma_{2,j}}\leq C_{2}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr)}.$ (E.5)
Furthermore, as $\sigma_{2}\leq\sigma_{m}\leq b_{2}$, the union bound and $\mathbb{P}\mathinner{(|\bm{z}|>t)}\leq 2\exp(-t^{2}/2)$ for $t\geq 0$ and $\bm{z}\sim\mathsf{N}_{1}(0,1)$ yields
@@
it follows that there exists a constant $K_{1}$ depending only on $b_{1},b_{2},c,$ and $m$ such that
@@
where the value of $K_{1}$ is suitably adjusted to justify the second inequality. Hence, by (E.4)–(E.6) there exists a constant $C_{3}$ depending only on $b_{1},b_{2},c$, and $m$ such that
$\mathbb{P}\mathinner{\biggl(A_{n}\leq C_{3}\sqrt{\log(dn)}\mathinner{\Bigl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\Bigr]}\biggr)}\geq 1-\frac{24}{n}-K_{1}r_{n}\geq 1-Kr_{n},$ (E.7)
where $K$ depends only on $b_{1},b_{2},c$, and $m$. Thus, writing
$\overline{A}_{n}\mathrel{\mathop{\ordinarycolon}}=C_{3}\sqrt{\log(dn)}\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]},$
@@
such that by Theorem 2.1 and the covariance matrix of $Y_{1}$ being $\Sigma_{0}=D^{-1}\Sigma D^{-1}$, there exists a constant $C$ depending only on $b_{1},b_{2},c$, and $m$ such that (using that $H^{\overline{A}_{n},\infty}\in\mathcal{H}$)
@@
$-(C+K)r_{n}-7\overline{A}_{n}\sqrt{\log(d)}\leq\mathbb{P}\mathinner{\bigl(S_{n,W,S}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z^{\prime}\in H\bigr)},$
@@
First, $\rho_{n,W,S}\leq\mathfrak{B}_{n}$, where $\mathfrak{B}_{n}$ is the upper bound on $\rho_{n,W,S}$ in Theorem 4.1. Next, by the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 of Chernozhuokov et al. (2022) (cf. also Proposition 2 in Chernozhukov et al. (2023a)),

### v4

@@ Appendix D Proofs for Section 3
$A_{n,j,k}\mathrel{\mathop{\ordinarycolon}}=\mathinner{\Bigl\{i\in\mathinner{\bigl\{1,\ldots,n\bigr\}}\mathrel{\mathop{\ordinarycolon}}\check{\alpha}_{j,k}\leq\tilde{U}_{i,j,k}\leq\check{\beta}_{j,k}\text{ and }\check{a}_{j}\leq\tilde{Y}_{i,j}\leq\check{b}_{j}\text{ and }\check{a}_{k}\leq\tilde{Y}_{i,k}\leq\check{b}_{k}\Bigr\}},$
one has that
$\phi_{\check{\alpha}_{j,k},\check{\beta}_{j,k}}\mathinner{(\tilde{U}_{i,j,k})}-\phi_{\check{a}_{j},\check{b}_{j}}(\tilde{Y}_{i,j})\phi_{\check{a}_{k},\check{b}_{k}}(\tilde{Y}_{i,k})=\tilde{U}_{i,j,k}-\tilde{Y}_{i,j}\tilde{Y}_{i,k}=0\qquad\text{for }i\in A_{n,j,k}.$
Hence,
$\mathfrak{S}_{n}$ $=\frac{1}{n}\sum_{i\in A_{n,j,k}^{c}}\mathinner{\bigl[\phi_{\check{\alpha}_{j,k},\check{\beta}_{j,k}}\mathinner{(\tilde{U}_{i,j,k})}-\phi_{\check{a}_{j},\check{b}_{j}}(\tilde{Y}_{i,j})\phi_{\check{a}_{k},\check{b}_{k}}(\tilde{Y}_{i,k})\bigr]},$
and for every $i\in A_{n,j,k}^{c}$
$\mathinner{\!\Bigl\lvert\phi_{\check{\alpha}_{j,k},\check{\beta}_{j,k}}\mathinner{(\tilde{U}_{i,j,k})}-\phi_{\check{a}_{j},\hat{b}_{j}}(\tilde{Y}_{i,j})\phi_{\check{a}_{k},\hat{b}_{k}}(\tilde{Y}_{i,k})\Bigr\rvert}\leq\mathinner{\bigl(|\check{\alpha}_{j,k}|\vee|\check{\beta}_{j,k}|\bigr)}+\mathinner{\bigl(|\check{a}_{j}|\vee|\check{b}_{j}|\bigr)}\mathinner{\bigl(|\check{a}_{k}|\vee|\check{b}_{k}|\bigr)}.$ (D.4)
To bound the right-hand side of (D.4), note that by Lemma G.7
$Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(U_{1,j,k})\leq\hat{\alpha}_{j,k}\leq\hat{\beta}_{j,k}\leq Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(U_{1,j,k})$
with probability at least $1-\frac{2}{d^{2}n}$, and where the second inequality used that $\varepsilon_{n}^{\prime}\in(0,1/2)$ by (9). Therefore, with at least this probability,
$|\check{\alpha}_{j,k}|\vee|\check{\beta}_{j,k}|\leq|Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(U_{1,j,k})|\vee|Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(U_{1,j,k})|$
$\mathinner{\bigl(|\check{a}_{j}|\vee|\check{b}_{j}|\bigr)}\mathinner{\bigl(|\check{a}_{k}|\vee|\check{b}_{k}|\bigr)}\leq$ $\mathinner{\bigl(|Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,j})|\vee|Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,j})|\bigr)}$
$\cdot$ $\mathinner{\bigl(|Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,k})|\vee|Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,k})|\bigr)}.$
$|Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(U_{1,j,k})|\vee|Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(U_{1,j,k})|+$ $\mathinner{\bigl(|Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,j})|\vee|Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,j})|\bigr)}$
$\cdot$ $\mathinner{\bigl(|Q_{c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,k})|\vee|Q_{1-c_{1,n}^{\prime}\varepsilon_{n}^{\prime}}(Y_{1,k})|\bigr)}.$
Next, recall that $\mathbb{E}U_{1,j,k}=\Sigma_{j,k}$$\mathbb{E}\mathinner{\lvert U_{1,j,k}-\mathbb{E}U_{1,j,k}\rvert}^{m/2}\leq 2^{m/2}\mathbb{E}|U_{1,j,k}|^{m/2}\leq 2^{m/2}\sigma_{m}^{m}$, $\mathbb{E}Y_{1,j}=0$, and $\mathbb{E}|Y_{1,j}|^{m}\leq\sigma_{m}^{m}$ for all $1\leq j,k\leq d$. Hence, Lemma G.1 implies that the previous display is bounded from above by
$\mathinner{\biggl(|\Sigma_{j,k}|+\frac{2\sigma_{m}^{2}}{(c_{1,n}^{\prime}\varepsilon_{n}^{\prime})^{2/m}}\biggr)}+\frac{\sigma_{m}^{2}}{(c_{1,n}^{\prime}\varepsilon_{n}^{\prime})^{2/m}}\leq\frac{4\sigma_{m}^{2}}{(c_{1,n}^{\prime}\varepsilon_{n}^{\prime})^{2/m}},$
the inequality following from $|\Sigma_{j,k}|\leq\sigma_{2}^{2}\leq\sigma_{m}^{2}$. Thus, since $|A_{n,j,k}^{c}|\leq 6\varepsilon_{n}^{\prime}n$, with probability at least $1-\frac{6}{d^{2}n}$ it holds that
@@
$\mathinner{\!\bigl\lvert\mathfrak{S}_{n,j,k}\bigr\rvert}\leq 6\varepsilon_{n}^{\prime}\cdot\frac{4\sigma_{m}^{2}}{(c_{1,n}^{\prime}\varepsilon_{n}^{\prime})^{2/m}}\leq\frac{24\sigma_{m}^{2}}{(c_{1,n}^{\prime})^{2/m}}\cdot{\varepsilon_{n}^{\prime}}^{1-\frac{2}{m}}\leq C_{3}\mathinner{\Bigl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(d^{2}n)}{n}\Bigr]}^{1-\frac{2}{m}}\Bigr)},$
where we inserted $\varepsilon_{n}^{\prime}$ from (7), used subadditivity of $z\mapsto z^{1-\frac{2}{m}}$, the lower bound on $c_{1,n}^{\prime}$ from Lemma G.6, and $C_{3}$ is the desired constant depending only on $b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$. Hence, (D.3) follows because $1-\frac{2}{m}\geq 1-\frac{1}{(m/2)\wedge 2}$ and $\frac{\log(d^{2}n)}{n}<1$ by (9).
@@
Note first that writing $\mathfrak{T}_{n}^{(3)}$ for the left-hand side of (9), it follows from $\frac{\log(d^{2}n)}{n}\leq{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}$ that
$\mathfrak{T}_{n}^{(3)}\leq 2\varepsilon_{n}^{\prime}+{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}+\sqrt{{\lambda_{2}^{\prime}}^{-2}{\varepsilon_{n}^{\prime}}^{2}+4{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}^{2}}\leq C_{1}\varepsilon_{n}^{\prime},$
for a constant $C_{1}$ depending only on $\lambda_{2}^{\prime}$.
Suppose first that $C_{1}\varepsilon_{n}^{\prime}<1$ such that (9) is satisfied by the previous display. By the triangle inequality $\tilde{\rho}_{n,W}$ is bounded from above by the sum of
@@
First, $\rho_{n,W}\leq\mathfrak{A}_{n}$, where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in Theorem 2.1. Next, by the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 of Chernozhukov et al. (2022) (cf. also Proposition 2 in Chernozhukov et al. (2023a)),
$B_{n}\leq C\mathinner{\Bigl(\log^{2}(d)\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}\Bigr)}^{1/2}\leq C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}^{1/2},$
the last inequality holding with probability at least $1-\frac{24}{n}$ by Theorem 3.1 and $C$ being a constant depending on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$ only (recall that (9) is satisfied).
If, on the other hand, $C_{1}\varepsilon_{n}^{\prime}\geq 1$ then by the definition of $\varepsilon_{n}^{\prime}$ in (7)
$\overline{\eta}_{n}\geq\frac{1}{2C_{1}\lambda_{1}^{\prime}}\qquad\text{or}\qquad\frac{\log(dn)}{n}=\frac{\log([dn]^{2})/2}{n}\geq\frac{\log(d^{2}n)}{2n}\geq\frac{1}{4C_{1}\lambda_{2}^{\prime}}.$
In either case, the upper bound on $\hat{\rho}_{n,W}$ remains valid by suitably adjusting $C$ (if needed).
That $\tilde{\rho}_{n,W}\to 0$ in probability follows from (B.7)–(B.9) of Lemma B.4. ∎
@@ Appendix E Proofs for Section 4
$\frac{\overline{\delta}}{\underline{\sigma}}\mathinner{\bigl(\sqrt{2\log(2d)}+2\bigr)}=\frac{\overline{\delta}}{\underline{\sigma}}\mathinner{\bigl(\sqrt{2(\log(2)+\log(d))}+2\bigr)}\leq\frac{\overline{\delta}}{\underline{\sigma}}\mathinner{\bigl(\sqrt{2\log(d)}+2+\sqrt{2(\log(2)}\bigr)},$
@@
We first establish (4.1). Writing $\mathfrak{T}_{n}^{(3)}$ for the left-hand side of (9), it follows from $\frac{\log(d^{2}n)}{n}\leq{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}$ that
$\mathfrak{T}_{n}^{(3)}\leq 2\varepsilon_{n}^{\prime}+{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}+\sqrt{{\lambda_{2}^{\prime}}^{-2}{\varepsilon_{n}^{\prime}}^{2}+4{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}^{2}}\leq K\varepsilon_{n}^{\prime},$
for a constant $K$ depending only on $\lambda_{2}^{\prime}$. Assume that
$K\varepsilon_{n}^{\prime}<1\qquad\text{and}\qquad C_{1}\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\leq\frac{b_{1}^{2}}{2},$ (E.3)
where $C_{1}$ is the constant from Theorem 3.1 depending only on $b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$. Note that $K\varepsilon_{n}^{\prime}<1$ implies, by the penultimate display, that (9) is satisfied which is used when invoking Theorem 3.1 below. If either of the conditions in (E.3) is violated, we conclude (4.1) with similar arguments as those employed in the end of the proofs of Theorems 2.1 and 3.2, respectively.
@@
$T_{n,j}=\frac{1}{\sqrt{n}\sigma_{2,j}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]},\qquad j=1,\ldots,d,$
@@
$\max_{j=1,\ldots,d}\mathinner{\!\biggl\lvert\frac{1}{\tilde{\sigma}_{n,j}}-\frac{1}{\sigma_{2,j}}\biggr\rvert}=\max_{j=1,\ldots,d}\frac{|\tilde{\sigma}_{n,j}-\sigma_{2,j}|}{\tilde{\sigma}_{n,j}\sigma_{2,j}}\leq C_{2}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}.$ (E.5)
Furthermore, as $\sigma_{2}\leq\sigma_{m}\leq b_{2}$, the union bound and $\mathbb{P}\mathinner{(|\bm{z}|>t)}\leq 2\exp(-t^{2}/2)$ for $t\geq 0$ and $\bm{z}\sim\mathsf{N}_{1}(0,1)$ yields for $Z\sim\mathsf{N}_{d}(0,\Sigma)$ that
@@
it follows that there exists a constant $K_{1}$ depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2},$ and $m$ such that
@@
where the value of $K_{1}$ is suitably adjusted to justify the second inequality. Hence, by (E.4)–(E.6) there exists a constant $C_{3}$ depending only on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$ such that
$\mathbb{P}\mathinner{\biggl(A_{n}\leq C_{3}\sqrt{\log(dn)}\mathinner{\Bigl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\Bigr]}\biggr)}\geq 1-\frac{24}{n}-K_{1}r_{n}\geq 1-Kr_{n},$ (E.7)
where $K$ depends only on $b_{1},b_{2},\lambda_{1},\lambda_{2}$, and $m$. Thus, writing
$\overline{A}_{n}\mathrel{\mathop{\ordinarycolon}}=C_{3}\sqrt{\log(dn)}\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]},$
@@
such that by Theorem 2.1 and the covariance matrix of $Y_{1}$ being $\Sigma_{0}=D^{-1}\Sigma D^{-1}$, there exists a constant $C$ depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2}$, and $m$ such that (using that $H^{\overline{A}_{n},\infty}\in\mathcal{H}$)
@@
$(C+K)r_{n}-7\overline{A}_{n}\sqrt{\log(d)}\leq\mathbb{P}\mathinner{\bigl(S_{n,W,S}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z^{\prime}\in H\bigr)},$
@@
First, $\rho_{n,W,S}\leq\mathfrak{B}_{n}$, where $\mathfrak{B}_{n}$ is the upper bound on $\rho_{n,W,S}$ in Theorem 4.1. Next, by the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 of Chernozhukov et al. (2022) (cf. also Proposition 2 in Chernozhukov et al. (2023a)),